ganeshie8 (ganeshie8):

find the integer solutions of \(px^2 + 2 = y^2\) where \(p\) is an odd prime

ganeshie8 (ganeshie8):

i found one by triall and error \(p=7\\ x=1\\ y=3\)

OpenStudy (jhannybean):

I have a feeling this is another proof ~.~ Haha.

OpenStudy (jhannybean):

Binomial? Fermats? The ones you've been working with?

ganeshie8 (ganeshie8):

i have no clue yet lol still trying...

ganeshie8 (ganeshie8):

wolfram rejects this straight away http://www.wolframalpha.com/input/?i=solve+px%5E2%2B2+%3D+y%5E2+over+integers

hartnn (hartnn):

x=1: p = y^2 -2 y=5, p =23 y= 7, p = 47 y=9, p =79 ... y must be odd in that case

ganeshie8 (ganeshie8):

Ahh thats interesting so it seems x = 1 produces infinitely many solutions as there are infinitely many primes of form y^2-2 thats neat :)

ganeshie8 (ganeshie8):

any easy way to conclude there are no solutions when \(x \ne 1\) ?

hartnn (hartnn):

\(y^2-2\) should not give out a prime when x not =1

hartnn (hartnn):

when divided by a perfect square*** how do we convert that into a mod equation? :O

hartnn (hartnn):

and why are we assuming there are no solutions when x not =1 there could be... ?

ganeshie8 (ganeshie8):

there is no easy way to work it wid mods i guess with squares...

ganeshie8 (ganeshie8):

yeah earlier i thought there can be only few solutions but after looking at ur reply it is clear that there are infinitely many solutions when x = 1 for x > 1 there could be solutions yeah.. idk i got this equation while attempting below problem which has only two solutions http://openstudy.com/study#/updates/54b12625e4b0dbc68afe844d looks i messed up somewhere hmm

ganeshie8 (ganeshie8):

im trying to find integer solutions of below equation : \[\frac{2^{p-1}-1}{p} = n^2\] which can be simplified like below wid some algebra \[\frac{2^{p-1}-1}{p} = \frac{ (2^{\frac{ p-1 }{ 2 }}-1)(2^{\frac{ p-1 }{ 2 }}+1) }{ p } = n^2\] \[ (2^{\frac{ p-1 }{ 2 }}-1)(2^{\frac{ p-1 }{ 2 }}+1) = pn^2\]

ganeshie8 (ganeshie8):

from there i got the equation we have been working with in this thread

ganeshie8 (ganeshie8):

\[ (2^{\frac{ p-1 }{ 2 }}-1)(2^{\frac{ p-1 }{ 2 }}+1) = pn^2\] since gcd of factors on left hand side is 1 we get below equations : \(2^{\frac{ p-1 }{ 2 }}-1 = px^2\) and \(2^{\frac{ p-1 }{ 2 }}+1 = y^2\) subtracting them i get \(px^2+2 = y^2\)

ganeshie8 (ganeshie8):

i expected to get only two solutions to this equation because the original equation \(\frac{2^{p-1}-1}{p} = n^2 \) has only two solutions.. idk whats wrong with my work above :/

OpenStudy (anonymous):

ur right , two solution according to Legendre theorem will work on this at home

OpenStudy (anonymous):

mostly Diophantine for quadratic could work

OpenStudy (anonymous):

for x=1 or -1 primes of the form y^2-2 works

OpenStudy (anonymous):

when x is not 1 y^2-2/x^2 is not integer this is what we need to show :O

OpenStudy (anonymous):

\(y=\pm 3,\pm 5,\pm 7,\pm 13,\pm 15,\pm 19.....\)

OpenStudy (anonymous):

interesting...

OpenStudy (anonymous):

>.<

ganeshie8 (ganeshie8):

hmm i see my silly mistake

OpenStudy (anonymous):

hmmm well i liked that only x=1

OpenStudy (anonymous):

great work in here

OpenStudy (kainui):

This reminds me of a problem I think either @No.name or @quickstudent posted a while back: p, q, and r are all primes, what are all the solutions to this equation: p^2-q^2 = r The method we figured out was that we could factor it to (p+q)(p-q) and since r is prime, that means one of these two has to be equal to 1, so that means p-q=1 since there's no way the other term can add two primes to get 1 haha. So then from this equation we notice p=1+q, but this implies they are consecutive primes! This only happens for q=2 and p=3. So we can check and see that 9-4 = 5 which is prime, so there is really only a single answer. I don't know, I'm kind of sick of Fermat and Goldbach right now lol.

ganeshie8 (ganeshie8):

I saw you and @No.name solving it :) I made a really silly mistake and ended up with this equation, i figured out the mistake very late... but it was fun trying to find the solutions of this equation xD

ganeshie8 (ganeshie8):

this is the error : http://gyazo.com/790f52d083ea95a54f5dade23a0f2d45 :'(

OpenStudy (kainui):

I might be on to something! Since we have that 2 there we know x and y are either both even or both odd. So we can represent them with this formula where a is either 0 or 1. \[\Large p(2n+a)^2+2 = (2m+a)^2\] AHA! Now look and see that if a=0 this implies both numbers are even. However this leads to a contradiction: \[\Large p(2n)^2 + 2 = (2m)^2 \\ ... \\ \Large 2pn^2 + 1 = 2m^2\] Obviously since both x and y have to be even or both odd, having them both even leads to only false statements that an odd number is equal to an even one! Now we know a=1 and we can plug that in...

OpenStudy (kainui):

I just did a deep burial and found some answers. =)

OpenStudy (kainui):

Let's start digging, shall we? We know that x and y are odd and we also know that p is odd, so p is of this form, p=2k+1, so I'll just selectively plug it in where p is added rather than multiplied since it won't matter since the bulk of the following arguments are around the idea that even*even and even*odd are both even, so we have to assume the added terms are even as well: \[p4(n)(n+1) + (p+1) = 4(m)(m+1) \\ p = 2k+1 \\ p4n(n+1) + 2k+2 = 4m(m+1) \\ p2n(n+1)+k+1=2m(m+1) \\ k=2h+1 \\ pn(n+1) + h+1 = m(m+1) \\ h = 2r+1\] This is where the reasoning stops, we know that h has to be odd because on either side we have two consecutive numbers multiplied together which means we have an even times an odd, so the number added (h+1) must have h odd so that it can give an even result. Substituting back up the hole we have: \[\LARGE p = 8r+7\]

OpenStudy (kainui):

@ganeshie8 @Marki I think you'll like this. =D

ganeshie8 (ganeshie8):

On cursory reading, that looks like an amazing finding XD the primes that satisfy this equation must be of form \(8k-1\) is it ? Im going trying to read the replies thoroughly now..

OpenStudy (kainui):

The primes must be of the form p=8r+7 where r is any integer, so we get p=7 when r=0 like you found. =D

OpenStudy (kainui):

r is any positive integer*

ganeshie8 (ganeshie8):

that means the primes cannot be of form 8k+1/8k+3/8k+5 thats really a neat result, im still going through the replies xD

OpenStudy (kainui):

That doesn't necessarily mean that they all work, I just showed they must be of this form. Upon doing the wolfram alpha test we find: http://www.wolframalpha.com/input/?i=15x%5E2%2B2%3Dy%5E2+integer+solutions

ganeshie8 (ganeshie8):

your first reply nicely concludes that x and y must be odd xD going thru second reply..

OpenStudy (anonymous):

Im very tired but i liked it XD

ganeshie8 (ganeshie8):

really cute @Kainui :) ill save the number theory tools to prove \(p \equiv -1 \pmod 8\) for now xD

ganeshie8 (ganeshie8):

Here is another way to get to that result w/o using so much of number theory : Any odd integer is of form \(4k\pm 1\) so the equation \(px^2 + 2 = y^2\) turns out to \[p(4m\pm1)^2 + 2 = (4n\pm 1)^2\] \[p(8x + 1) + 2 = (8y+ 1)^2\] dividing by \(8\) we get \[p(1) + 2 \equiv 1 \pmod {8}\] \[p \equiv -1\equiv 7\pmod {8}\] this is exactly same as your method haha!

ganeshie8 (ganeshie8):

if we want to use number theory sophistication, there is this pell's equation which im scared of messing wid..

ganeshie8 (ganeshie8):

http://mathworld.wolfram.com/PellEquation.html

OpenStudy (kainui):

Haha interesting! I am going to have to take a break, I am tired too. But I will leave one last cute bit of nonsense, we could also rewrite the consecutive terms as if it was a sum of numbers like this to divide out one more set of 2's. \[\Large n(n+1) = 2 S_n = 2 \sum_{k=1}^n k\]

ganeshie8 (ganeshie8):

you're messing wid triangular numbers haha

OpenStudy (kainui):

I prefer to call triangular numbers "baby gauss numbers" since they always remind me of that story how he added them all up lol.

ganeshie8 (ganeshie8):

lol anythign to do with sum of first n natural numbers remind me the same xD nth triangular can be given by below short form \[S_n = \binom{n+1}{2} \]

ganeshie8 (ganeshie8):

using "t" for triangular is more appropriate i feel \[t_n = \binom{n+1}{2} \]

ganeshie8 (ganeshie8):

\[t_{n-1} + t_n = n^2\] greeks were really excited about properties of these numbers

OpenStudy (kainui):

It's kind of obvious geometrically, so I can see how they would have discovered this since algebra kinda makes it look cooler than it really is haha.